OpenStudy (anonymous):

What is the sum of the arithmetic sequence 135, 123, 111.. if there are 32 terms?

OpenStudy (campbell_st):

\[s _{n} = \frac{n}{2}[2a + (n-1)d]\] in theis question 1st term a = 135, number of terms n = 135 and common difference d= -12 substitute them into the formula and evaluate

OpenStudy (campbell_st):

oops n = 32

OpenStudy (anonymous):

I am confused on what to do.

OpenStudy (anonymous):

\[a _{n} = a _{1} + (n - 1)d\] \[a _{n} = 135 + (32 - 1)-12\] \[a _{n} = 135 + (31)(-12)\] \[a _{n} = 135 + (31)(-12)\] \[a _{n} = 135 - 372\] \[a _{n} = -237\] \[S_{n} = n/2(a_{1} + a_{n})\] \[S_{n} = 32/2(135 - 237)\] \[S_{n} = 16(-102)\] \[S_{n} = -1632\]

OpenStudy (campbell_st):

both methods shown use the idea of the sum of the 1st and last term multiplied by the number of terms over 2 will give the sum do you know how to find a term in an arithmetic series...?

OpenStudy (anonymous):

No.

OpenStudy (campbell_st):

so you know any formula for an arithmetic series...?

OpenStudy (anonymous):

no. Can you please just give me the answer? Because I have to be done with this in 30 minutes to graduate.